OpenStudy (anonymous):

What is the area of a circle with a diameter of fourteen meters? Use twenty-two sevenths for pi.

OpenStudy (anonymous):

A=pi*r² I didn't understand this "use twenty-two sevenths for pi" thing.

OpenStudy (anonymous):

Here are the choices it gives me.. one hundred thirty-two square meters one hundred fifty-four square meters one hundred seventy-six square meters one hundred ninety-eight square meters

OpenStudy (anonymous):

A = 3,14*7² = 154m²